3. The first excited state of Ca is reached by absorption of 422.7 nm light. What is the energy difference between the ground state and this excited state?
Answer #1

Energy difference (E) = hc/l


   h = planks constant = 6.625*10^-34 j.s

   c = light velocity = 3*10^8 m/s

   l = wavelength = 422.7 nm = 4.227*10^-7 m

       = (6.625*10^-34*3*10^8/(4.227*10^-7))

   = 4.7*10^-19 joule.